OpenStudy (anonymous):

[10.02] Solve the proportion 2 times x all over 7 times x minus 6 equals 2 over x.. What is the value of x? x = -1 and x = -7 x = 1 and x = -7 x = 1 and x = 6 x = -1 and x = 6

OpenStudy (anonymous):

\[\frac{ 2x }{ 7x-6 }=\frac{ 2 }{ x }\] is that it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so cross multiply so: \[2(7x-6)=2x*x\] so:\[2x^2=14x-12\] rewrite that as: \[2x^2-14x+12=0\] what factors of 2*12=24 equal -14 so -12 and -2 multiply to 24 and add up to -14 so write the equation as:\[2x^2-12x-2x+12=0\] factor by grouping so: \[2x(x-6)-2(x-6)=0\] factor: \[(2x-2)(x-6)=0\] so: \[2x-2=0\] or \[x-6=0\] so: \[x=6\]or x=1

OpenStudy (anonymous):

the answer is x=1 and x=6

OpenStudy (anonymous):

oh ok i got it now thank you